# Robust mass lumping and outlier removal strategies in isogeometric analysis

### v1

@@ Abstract
Mass lumping techniques are commonly employed in explicit time integration schemes for problems in structural dynamics and both avoid solving costly linear systems with the consistent mass matrix and increase the critical time step. In isogeometric analysis, the critical time step is constrained by so-called “outlier” frequencies, representing the inaccurate high frequency part of the spectrum. Removing or dampening these high frequencies is paramount for fast explicit solution techniques. In this work, we propose robust mass lumping and outlier removal techniques for nontrivial geometries, including multipatch and trimmed geometries. Our lumping strategies provably do not deteriorate (and often improve) the CFL condition of the original problem and are combined with deflation techniques to remove persistent outlier frequencies. Numerical experiments reveal the advantages of the method, especially for simulations covering large time spans where they may halve the number of iterations with little or no effect on the numerical solution.
@@ 1 Introduction and background
Isogeometric analysis is a discretization technique for solving partial differential equations (PDEs) which relies on spline functions such as B-splines and non-uniform rational B-splines (NURBS) both for parameterizing the geometry and representing the solution [1, 2]. Spline functions used in isogeometric analysis offer distinctive advantages over Lagrange polynomials used in classical finite element discretizations, including exact representation of common geometries and superior approximation properties [3, 4, 5]. In structural dynamics, the advantages of isogeometric analysis were already evidenced in [6, 7, 8] where maximally smooth spline approximations removed the so-called “optical branches” from the discrete spectrum, a typical artifact of classical finite element discretizations. As a matter of fact, nearly all discrete eigenvalues approximate the continuous eigenvalues with high accuracy except for the largest ones. They form noticeable spikes in the upper part of the spectrum and were coined “outlier” eigenvalues for this very reason [6]. These inaccurate eigenvalues grow with the mesh size and spline order [9] thereby severely constraining the critical time step of explicit time integration schemes, often referred to as the Courant–Friedrichs–Lewy (CFL) condition. For example, for undamped dynamical systems, the critical time step of the central difference method is
@@
Since the advent of isogeometric analysis, much effort has focused on removing the outliers from the discrete spectrum. A nonlinear spline parametrization was first proposed in [6] by uniformly distributing control points by changing the geometry parametrization but defeats the spirit of isogeometric analysis. The method also reportedly undermines the accuracy of the low frequencies and modes [12]. A similar method was later proposed in [13] by constructing “smoothed” knot vectors that are approximations to optimal knot vectors (related to some $n$-width optimal spline spaces) but suffer from similar drawbacks. In [14] the authors proposed to instead use the $n$-width optimal spline spaces in [15] to remove outliers, and these spaces were later proven to be outlier-free in [16] without loss of accuracy in the low frequencies. These optimal spaces mimic the true eigenfunctions by imposing certain higher-order derivatives to be zero at the boundary. Bases for these spaces were constructed in [17, 15, 18, 16] by using certain symmetry properties of B-splines. A similar strategy was numerically observed to be outlier-free in [12] for the Laplacian, however, some outliers were still observed for the biharmonic problem with the spline spaces proposed in [12]. By mimicking the properties of the $n$-width optimal spline spaces in [19] outliers for the biharmonic problem were later completely removed in [20]. In both [12] and [20] bases for the considered spline spaces were computed using a similar strategy as the MDB-spline extraction technique developed in [21]. A related strategy to the above was devised in [22] where the authors weakly impose the constraints coming from the true eigenfunctions by penalizing high order derivatives near the boundary. In the multipatch setting such a strategy was also used in [23] by weakly enforcing $C^{p-1}$ continuity at patch interfaces to remove interior outliers. These penalization techniques often take the form of mass scaling, which is also widely employed for classical finite element methods [24, 25, 26]. Rather than completely removing outliers, they strongly mitigate them but also require heuristics or dedicated algorithms for computing the generally unknown penalization parameters. We note that these approaches could be combined to mitigate both boundary and interior outliers resulting from $C^{0}$ continuity at patch interfaces. Also in the context of classical finite element methods, an eigenvalue deflation technique was proposed in [27, 28] by explicitly computing the largest eigenvalues and eigenmodes. However, the authors do not discuss the computational overhead in their experiments and applying their method to the global assembled mass matrix is generally infeasible due to the large number of inaccurate high frequencies in classical $C^{0}$ finite element methods.
@@
For applications in explicit dynamics, restrictive CFL conditions are not the only issue. Due to the inherent cost of “exactly” solving linear systems with the mass matrix [29, 30], obtaining an easily invertible (preferably diagonal or tridiagonal) mass matrix is paramount. Mass lumping has historically been used for approximating the consistent mass matrix by a diagonal (lumped) mass matrix. Among the scores of methods proposed in the 1970s for classical finite element methods, only a handful are applicable to isogeometric analysis due to the non-interpolatory nature of the basis functions. Among them is the classical row-sum technique [31, 10]. Provably, the row-sum technique does not worsen the CFL condition of the consistent mass [32]. However, for isogeometric analysis, the method performs poorly in higher dimensions and strong numerical evidence suggests it reduces the converge rate to quadratic order, independently of the spline order [6]; a property only proved for 1D problems and low spline orders [6]. Since then much research effort has focused on devising more accurate and potentially high order mass lumping schemes for isogeometric analysis. Cottrell et al. [2] first suggested constructing diagonal mass matrices by using dual basis functions as test functions in a Petrov-Galerkin framework. However, computer implementations come with all sorts of difficulties and initially the idea did not gain much momentum until it was taken up again in [33] with promising results. Since then, there has been a surge of interest. In [34], high order convergence is achieved by combining (approximate) dual basis functions with the row-sum technique. However, the implementation remains technical and ongoing research is focusing on alleviating these issues, in particular related to the imposition of boundary conditions [35]. In another line of research, families of banded and Kronecker product matrices were constructed in [32] by increasing the bandwidth of the row-sum lumped mass matrix and were shown to significantly improve the accuracy. Unfortunately, such improvements are only realized on trivial geometries and are tied to an improvement of the constant instead of the convergence rate.
@@
Mass lumping may sometimes dramatically impact the CFL condition. For trimmed geometries [36], Leidinger [37] first showed that the CFL condition was not affected by small trimmed elements if the mass matrix was lumped. This finding was further supported by the studies in [38, 39] but also raised concerns over the accuracy of the smallest eigenvalues and modes. Devising accurate mass lumping/scaling techniques for trimmed geometries is an ongoing challenge.
@@
Fast solution methods in explicit dynamics usually combine outlier removal, mass scaling and mass lumping in a sometimes ad hoc fashion. In this article, we propose such a strategy for nontrivial geometries and provide a strong mathematical foundation for our method. The article is structured as follows: in Section 3, we first recall the mass lumping techniques devised in [32] and then design new mass lumping techniques for nontrivial single-patch and multipatch problems. Although these techniques provably do not worsen the CFL condition of the original problem, they may not significantly improve it either. Thus, they are combined in Section 4 with outlier removal techniques that deflate the spectrum from persistent outlier eigenvalues and generalize the method proposed in [27, 28]. Contrary to classical $C^{0}$ finite elements, isogeometric discretizations typically feature a small number of rapidly increasing eigenvalues towards the end of the spectrum, for which deflation techniques are well-suited. Section 5 gathers some numerical experiments illustrating the theoretical findings and demonstrating the advantages of the method. Finally, conclusions are drawn in Section 6.
@@ 2 Model problem and its discretization
where $u_{0}$ and $v_{0}$ are some initial displacement and velocity, respectively, $\rho$ and $\kappa$ are some positive valued coefficient functions and we prescribe homogeneous Dirichlet boundary conditions for simplicity. In a standard Galerkin discretization, we look for an approximation $u_{h}(.,t)$ of $u(.,t)$ in a finite dimensional subspace $V_{h}$ and test against all functions in $V_{h}$, which leads to solving the semi-discrete problem (see for instance [10, 40])
@@
where $K,M\in\mathbb{R}^{n\times n}$ are the stiffness and mass matrices, respectively. The time-dependent right-hand side vector $\mathbf{f}(t)\in\mathbb{R}^{n}$ accounts for the function $f$ and potential non-homogeneous Neumann and Dirichlet boundary conditions. Finally, $\mathbf{u}(t)\in\mathbb{R}^{n}$ is the coefficient vector of the approximate solution $u_{h}(\mathbf{x},t)$ in a basis of $V_{h}$. Isogeometric analysis consists in choosing spline functions from computer-aided-design (CAD) such as B-splines both for representing the approximate solution and describing the geometry [1, 2]. Such functions follow a standardized construction in a so-called parametric domain $\hat{\Omega}=(0,1)^{d}$ before being defined in the physical domain $\Omega$. In dimension $d=1$, the B-spline basis $\{\hat{B}_{i}\}_{i=1}^{n}$ is constructed recursively from a knot vector $\Xi:=(\xi_{1},\dots,\xi_{n+p+1})$, which is a sequence of non-decreasing real values. The integers $p$ and $n$ denote the spline degree and spline space dimension, respectively. A knot vector is called open if
@@
Internal knots of multiplicity $1\leq m\leq p$ give rise to $C^{k}$ continuous spline spaces, denoted $\mathcal{S}^{k}_{p,\Xi}$, where $k=p-m$. Our work primarily focuses on (but is not restricted to) maximally smooth $C^{p-1}$ spaces obtained when the multiplicity of each internal knot is $1$; i.e. the so-called isogeometric $k$-method. In dimension $d\geq 2$, the spline space is defined as a tensor product of univariate spaces, which all follow a similar construction. The degree, space dimension and continuity along each direction are collected in the vectors $\mathbf{p}=(p_{1},p_{2},\dots,p_{d})$, $\mathbf{n}=(n_{1},n_{2},\dots,n_{d})$ and $\mathbf{k}=(k_{1},k_{2},\dots,k_{d})$, respectively, and we denote the resulting spline space $\mathcal{S}^{\mathbf{k}}_{\mathbf{p},\mathbf{\Xi}}$ (where the dependency on the knot vectors $\Xi_{1},\dots,\Xi_{d}$ is specified by $\mathbf{\Xi}$). In dimension $d\geq 2$, it becomes convenient to label basis functions with multi-indices $\mathbf{i}=(i_{1},i_{2},\dots,i_{d})$ which are often identified with “linear” indices in the global numbering. This identification permits a slight abuse of notation when writing
@@
$K_{ij}=\int_{\hat{\Omega}}(\nabla\hat{B}_{i}(\hat{\mathbf{x}}))^{T}G(\hat{\mathbf{x}})\nabla\hat{B}_{j}(\hat{\mathbf{x}})\quad\text{and}\quad M_{ij}=\int_{\hat{\Omega}}c(\hat{\mathbf{x}})\hat{B}_{i}(\hat{\mathbf{x}})\hat{B}_{j}(\hat{\mathbf{x}})\qquad 1\leq i,j\leq n$ (2.3)
@@
For discretizing (2.2) in time, explicit methods are usually preferred for fast dynamic processes such as blasts or impacts due to the physical restriction on the step size. Scores of methods have been proposed in the literature, including the central difference, Wilson-$\theta$ and generalized $\alpha$ methods to name just a few. Most of them are commonly included in textbooks [10, 11], which the reader may consult for details. The critical time step depends on the method (see e.g. (1.1) for the central difference method in the undamped case) but in the undamped case all explicit methods require solving a linear system with the mass matrix at least once in each iteration and is the reason for mass lumping, which we describe in the next section.
@@ 3.1 Row-sum mass lumping and its generalization
Despite the apparent downgrade in convergence rate, the row-sum technique remains very popular owing to its simplicity and straightforward implementation. We define it through the application of a lumping operator [32].
@@
The Loewner partial order is a natural choice in our context since it is the matrix equivalent of bounding bilinear forms. Indeed, if $M$ and $\tilde{M}$ are some matrices stemming from the symmetric bilinear forms $b,\tilde{b}\colon V_{h}\times V_{h}\to\mathbb{R}$, respectively, then
@@
When the construction of $\tilde{M}$ is completely algebraic, we usually do not have an explicit representation of $\tilde{b}(u_{h},u_{h})$, but we might still bound it by understanding the relation between $\tilde{M}$ and $M$ in the Loewner ordering. The argument has already been used to good effect in [32] where it was shown that for two symmetric positive definite matrices $A,B\in\mathbb{R}^{n\times n}$, the generalized eigenvalues of $(A,\mathcal{L}(B))$ are always smaller or equal to those of $(A,B)$, which is a consequence of the fact that $\mathcal{L}(B)\succeq B$ [32, Corollary 3.10]. The authors in [32] used eigenvalue bounds to prove the result but there exist multiple short and elegant ways of reaching the same conclusion. For instance, if $B$ is nonnegative and one thinks of it as a weighted adjacency matrix, then $\mathcal{L}(B)$ is the degree matrix and $\mathcal{L}(B)-B$ is its graph Laplacian, which plays a prominent role in spectral clustering techniques [43]. A direct computation then shows that $\mathbf{x}^{T}(\mathcal{L}(B)-B)\mathbf{x}=\frac{1}{2}\sum_{i,j=1}^{n}(x_{i}-x_{j})^{2}b_{ij}\geq 0$ for any $\mathbf{x}\in\mathbb{R}^{n}$. This technique will be employed later in our paper.

### v2

# Mass lumping and outlier removal strategies for complex geometries in isogeometric analysis

@@ Abstract
Mass lumping techniques are commonly employed in explicit time integration schemes for problems in structural dynamics and both avoid solving costly linear systems with the consistent mass matrix and increase the critical time step. In isogeometric analysis, the critical time step is constrained by so-called “outlier” frequencies, representing the inaccurate high frequency part of the spectrum. Removing or dampening these high frequencies is paramount for fast explicit solution techniques. In this work, we propose mass lumping and outlier removal techniques for nontrivial geometries, including multipatch and trimmed geometries. Our lumping strategies provably do not deteriorate (and often improve) the CFL condition of the original problem and are combined with deflation techniques to remove persistent outlier frequencies. Numerical experiments reveal the advantages of the method, especially for simulations covering large time spans where they may halve the number of iterations with little or no effect on the numerical solution.
@@ 1 Introduction and background
Isogeometric analysis is a discretization technique for solving partial differential equations (PDEs) that relies on spline functions such as B-splines and non-uniform rational B-splines (NURBS) both for parameterizing the geometry and representing the solution [1, 2]. Spline functions used in isogeometric analysis offer distinctive advantages over Lagrange polynomials used in classical finite element discretizations, including exact representation of common geometries and superior approximation properties [3, 4, 5]. In structural dynamics, the advantages of isogeometric analysis were already evidenced in [6, 7, 8] where maximally smooth spline approximations removed the so-called “optical branches” from the discrete spectrum, a typical artifact of classical finite element discretizations. As a matter of fact, nearly all discrete eigenvalues approximate the continuous eigenvalues with high accuracy except for the largest ones. They form noticeable spikes in the upper part of the spectrum and were coined “outlier” eigenvalues for this very reason [6]. These inaccurate eigenvalues grow with the mesh size and spline order [9] thereby severely constraining the critical time step of explicit time integration schemes, often referred to as the Courant–Friedrichs–Lewy (CFL) condition. For example, for undamped dynamical systems, the critical time step of the central difference method is
@@
Since the advent of isogeometric analysis, much effort has focused on removing the outliers from the discrete spectrum. A nonlinear spline parametrization was first proposed in [6] by uniformly distributing control points and changing the geometry parametrization but defeats the spirit of isogeometric analysis. The method also reportedly undermines the accuracy of the low frequencies and modes [12]. A similar strategy was later proposed in [13] by constructing “smoothed” knot vectors that are approximations to optimal knot vectors (related to some $n$-width optimal spline spaces) but suffers from similar drawbacks. Instead, in [14], the authors used the $n$-width optimal spline spaces defined in [15] to remove outliers, and these spaces were later proven to be outlier-free in [16] without loss of accuracy in the low frequencies. These optimal spaces mimic the true eigenfunctions by imposing certain higher-order derivatives to be zero at the boundary. Bases for these spaces were constructed in [17, 15, 18, 16] by using certain symmetry properties of B-splines. A similar strategy was numerically observed to be outlier-free in [12] for the Laplacian, however, some outliers were still observed for the biharmonic problem with the spline spaces proposed in [12]. By mimicking the properties of the $n$-width optimal spline spaces in [19], outliers for the biharmonic problem were later completely removed in [20]. In both [12] and [20], the basis construction resembles the MDB-spline extraction technique developed in [21]. A related outlier removal technique was devised in [22], where the authors weakly impose the constraints coming from the true eigenfunctions by penalizing high order derivatives near the boundary. This same strategy was later employed in [23] in the multipatch setting by weakly enforcing $C^{p-1}$ continuity at patch interfaces to remove interior outliers. These penalization techniques often take the form of mass scaling, which is also widely employed for classical finite element methods [24, 25, 26]. Rather than completely removing outliers, they strongly mitigate them but also require heuristics or dedicated algorithms for computing the generally unknown penalization parameters. We note that these approaches could be combined to mitigate both boundary and interior outliers resulting from $C^{0}$ continuity at patch interfaces. Also in the context of classical finite element methods, an eigenvalue deflation technique was proposed in [27, 28] by explicitly computing the largest eigenvalues and eigenmodes. However, the authors do not discuss the computational overhead in their experiments and applying their method to the global assembled mass matrix is generally infeasible due to the large number of inaccurate high frequencies in classical $C^{0}$ finite element methods.
@@
For applications in explicit dynamics, restrictive CFL conditions are not the only issue. Due to the inherent cost of “exactly” solving linear systems with the mass matrix [29, 30], obtaining an easily invertible (preferably diagonal or tridiagonal) mass matrix is paramount. Mass lumping has historically been used for approximating the consistent mass matrix by a diagonal (lumped) mass matrix. Among the scores of methods proposed in the 1970s for classical finite element methods, only a handful are applicable to isogeometric analysis due to the non-interpolatory nature of the basis functions. Among them is the classical row-sum technique [31, 10]. Provably, the row-sum technique does not worsen the CFL condition of the consistent mass [32]. However, for isogeometric analysis, the method performs poorly in higher dimensions and strong numerical evidence suggests it reduces the converge rate to quadratic order, independently of the spline order [6]; a property only proved for 1D problems and low spline orders [6]. Since then, much research effort has focused on devising more accurate and potentially high order mass lumping schemes for isogeometric analysis. Cottrell et al. [2] first suggested constructing diagonal mass matrices by using dual basis functions as test functions in a Petrov-Galerkin framework. However, computer implementations come with all sorts of difficulties and initially the idea did not gain much momentum until it was taken up again in [33] with promising results, sparking a surge of interest in dual bases. In [34], high order convergence is achieved by combining (approximate) dual basis functions with the row-sum technique. Unfortunately, the method is currently limited to single-patch geometries and extensions to multipatch or trimmed geometries are not in sight. Even in the single-patch case, several technical difficulties still hold it back, such as the imposition of boundary conditions, which is the focus of ongoing research [35]. In another line of research, families of banded and Kronecker product matrices were constructed in [32] by increasing the bandwidth of the row-sum lumped mass matrix and were shown to significantly improve the accuracy. Unfortunately, such improvements are only realized on trivial geometries and are tied to an improvement of the constant instead of the convergence rate.
@@
Mass lumping may sometimes dramatically impact the CFL condition. For trimmed geometries [36], Leidinger [37] first showed that the CFL condition was not affected by small trimmed elements if the mass matrix was lumped provided the spline’s order and smoothness were sufficiently large with respect to the order of the differential operator. This finding was further supported by the studies in [38, 39, 40] but also raised concerns over the accuracy of the smallest eigenvalues and modes. Devising accurate mass lumping/scaling techniques for trimmed geometries is an ongoing challenge.
@@
Fast solution methods in explicit dynamics usually combine outlier removal, mass scaling and mass lumping in a sometimes ad hoc fashion. In this article, we propose such a strategy for complex geometries in isogeometric analysis and is substantiated by a strong mathematical foundation. Our contributions are twofold: firstly, we extend the mass lumping techniques proposed in [32] to nontrivial single-patch, multipatch and trimmed geometries and secondly propose an algebraic outlier removal strategy compatible with any lumping technique. Although this method has been independently proposed for classical finite element discretizations [27, 28], we generalize it and thoroughly discuss its implementation and computational cost. The article is structured as follows: after describing the model problem and its discretization in Section 2, we recall in Section 3 the mass lumping techniques devised in [32] and later extend them to nontrivial single-patch and multipatch problems. Although these techniques provably do not worsen the CFL condition of the original problem, they may not significantly improve it either. Thus, they are combined in Section 4 with outlier removal techniques that deflate the spectrum from persistent outlier eigenvalues and generalize the method proposed in [27, 28]. Contrary to classical $C^{0}$ finite elements, isogeometric discretizations typically feature a small number of rapidly increasing eigenvalues towards the end of the spectrum, for which deflation techniques are well-suited. Section 5 gathers some numerical experiments illustrating the theoretical findings and demonstrating the advantages of the method. Finally, conclusions are drawn in Section 6.
@@ 2 Model problem and its discretization
where $u_{0}$ and $v_{0}$ are some initial displacement and velocity, respectively, $\rho$ and $\kappa$ are some positive-valued coefficient functions and we prescribe homogeneous Dirichlet boundary conditions for simplicity. In a standard Galerkin discretization, we look for an approximation $u_{h}(.,t)$ of $u(.,t)$ in a finite dimensional subspace $V_{h}$ and test against all functions in $V_{h}$, which leads to solving the semi-discrete problem (see for instance [10, 41])
@@
where $K,M\in\mathbb{R}^{n\times n}$ are the stiffness and mass matrices, respectively. The time-dependent right-hand side vector $\mathbf{f}(t)\in\mathbb{R}^{n}$ accounts for the function $f$ and potential non-homogeneous Neumann and Dirichlet boundary conditions. Finally, $\mathbf{u}(t)\in\mathbb{R}^{n}$ is the coefficient vector of the approximate solution $u_{h}(\mathbf{x},t)$ in a basis of $V_{h}$. Isogeometric analysis consists in choosing spline functions from computer-aided-design (CAD) such as B-splines both for representing the approximate solution and describing the geometry [1, 2]. Such functions follow a standardized construction in a so-called parametric domain $\hat{\Omega}=(0,1)^{d}$ before being defined in the physical domain $\Omega$. In dimension $d=1$, the B-spline basis $\{\hat{B}_{i}\}_{i=1}^{n}$ is constructed recursively from a knot vector $\Xi:=(\xi_{1},\dots,\xi_{n+p+1})$, which is a sequence of non-decreasing real numbers. The integers $p$ and $n$ denote the spline degree and spline space dimension, respectively. A knot vector is called open if
@@
Internal knots of multiplicity $1\leq m\leq p$ give rise to $C^{k}$ continuous spline spaces, denoted $\mathcal{S}^{k}_{p,\Xi}$, where $k=p-m$. Our work primarily focuses on (but is not restricted to) maximally smooth $C^{p-1}$ spaces obtained when the multiplicity of each internal knot is $1$; i.e. the so-called isogeometric $k$-method. In dimension $d\geq 2$, the spline space is defined as a tensor product of univariate spaces, which all follow a similar construction. The degree, space dimension and continuity along each direction are collected in the vectors $\mathbf{p}=(p_{1},p_{2},\dots,p_{d})$, $\mathbf{n}=(n_{1},n_{2},\dots,n_{d})$ and $\mathbf{k}=(k_{1},k_{2},\dots,k_{d})$, respectively, and we denote the resulting spline space $\mathcal{S}^{\mathbf{k}}_{\mathbf{p},\mathbf{\Xi}}$ (where the dependency on the knot vectors $\Xi_{1},\dots,\Xi_{d}$ is specified by $\mathbf{\Xi}$). In dimension $d\geq 2$, it becomes convenient to label basis functions with multi-indices $\mathbf{i}=(i_{1},i_{2},\dots,i_{d})$, which are often identified with “linear” indices in the global numbering. This identification permits a slight abuse of notation when writing
@@
$K_{ij}=\int_{\hat{\Omega}}(\nabla\hat{B}_{i}(\hat{\mathbf{x}}))^{T}G(\hat{\mathbf{x}})\nabla\hat{B}_{j}(\hat{\mathbf{x}})\quad\text{and}\quad M_{ij}=\int_{\hat{\Omega}}c(\hat{\mathbf{x}})\hat{B}_{i}(\hat{\mathbf{x}})\hat{B}_{j}(\hat{\mathbf{x}})\quad 1\leq i,j\leq n$ (2.3)
@@
For discretizing (2.2) in time, explicit methods are usually preferred for fast dynamic processes such as blasts or impacts due to the physical restriction on the step size. Scores of methods have been proposed in the literature, including the central difference, Wilson-$\theta$ and generalized $\alpha$ methods to name just a few. Most of them are commonly included in textbooks [10, 11], which the reader may consult for details. The critical time step depends on the method (see e.g. (1.1) for the central difference method in the undamped case) but in the undamped case all explicit methods require solving a linear system with the mass matrix at least once in each iteration. Although very efficient preconditioning strategies have been devised [44, 45, 13, 46], practitioners still mostly resort to mass lumping, which we describe in the next section.
@@ 3.1 Row-sum mass lumping and its generalization
As for preconditioning matrices, lumped mass matrices are easily invertible approximations of the consistent mass. However, contrary to preconditioning, mass lumping consists in directly substituting the consistent mass with its approximation and therefore, it should not only provide a sufficiently accurate approximation but also guarantee an improvement of the CFL condition. Several techniques have been developed but despite the apparent downgrade in convergence rate, the row-sum technique remains very popular owing to its simplicity and straightforward implementation. We define it through the application of a lumping operator [32, Definition 3.7].
@@
Remark 3.2.
Note that the absolute values in the above definition are unnecessary for isogeometric analysis due to the positivity of the B-spline basis. However, in the case of a more general basis, they ensure that $\mathcal{L}(M)\succeq M$ and subsequently guarantee an improvement of the CFL [32, Corollary 3.10]. This result generally does not hold without them, as is well-known in the finite element community. Of course, the accuracy of lumping with absolute values when the basis is allowed to be negative is a whole separate issue and should be studied on a case-by-case basis.
@@
The Loewner partial order is a natural choice in our context since it is the matrix equivalent of bounding bilinear forms. Indeed, if $M$ and $\tilde{M}$ are the Gram matrices for the symmetric bilinear forms $b,\tilde{b}\colon V_{h}\times V_{h}\to\mathbb{R}$ in some basis of $V_{h}$, then
@@
When the construction of $\tilde{M}$ is completely algebraic, we rarely have an explicit representation of $\tilde{b}(u_{h},u_{h})$, but we might still bound it by understanding the relation between $\tilde{M}$ and $M$ in the Loewner ordering. The argument has already been used to good effect in [32] where it was shown that for two symmetric positive definite matrices $A,B\in\mathbb{R}^{n\times n}$, the generalized eigenvalues of $(A,\mathcal{L}(B))$ are always smaller or equal to those of $(A,B)$, which is a consequence of the fact that $\mathcal{L}(B)\succeq B$ [32, Corollary 3.10]. The authors in [32] used eigenvalue bounds to prove the result but there exist multiple short and elegant ways of reaching the same conclusion. For instance, if $B$ is nonnegative and one thinks of it as a weighted adjacency matrix, then $\mathcal{L}(B)$ is the degree matrix and $\mathcal{L}(B)-B$ is its graph Laplacian, which plays a prominent role in spectral clustering techniques [47]. A direct computation then shows that $\mathbf{x}^{T}(\mathcal{L}(B)-B)\mathbf{x}=\frac{1}{2}\sum_{i,j=1}^{n}(x_{i}-x_{j})^{2}b_{ij}\geq 0$ for any $\mathbf{x}\in\mathbb{R}^{n}$. This technique will be employed later in our paper.